# Optimal Piecewise-based Mechanism for Collecting Bounded Numerical Data under Local Differential Privacy

### v1

@@ Abstract.
This paper investigates the optimal design of piecewise-based mechanisms to maximize data utility under LDP. We demonstrate that existing piecewise-based mechanisms are heuristic forms of the $3$-piecewise mechanism, which is far from enough to study optimality. We generalize the $3$-piecewise mechanism to its most general form, i.e. $m$-piecewise mechanism with no pre-defined form of each piece. Under this form, we derive the closed-form optimal mechanism by combining analytical proofs and off-the-shelf optimization solvers. Next, we extend the generalized piecewise-based mechanism to the circular domain (along with the classical domain), defined on a cyclic range where the distance between the two endpoints is zero. By incorporating this property, we design the optimal mechanism for the circular domain, achieving significantly improved data utility compared with existing mechanisms.
@@ 1. Introduction
Numerical data with bounded domains is a fundamental data type in personal devices. These bounded domains can be categorized into two types: linear ranges, such as sensor readings in $[0,1)$, referred to as classical domain; and cyclic ranges, such as angular measurements in $[0,2\pi)$, referred to as circular domain. These types of data are crucial for third-party platforms to provide personalized services. However, collecting such data involves particular privacy concerns, as the third-party collectors are potentially untrusted and the data often contain sensitive information. Simple anonymization techniques (Yang et al., 2013; Niu et al., 2014) have been proven insufficient to prevent privacy leakage (Wikipedia contributors, 2023; Narayanan and Shmatikov, 2008; Yu et al., 2007). Therefore, a provable privacy guarantee is necessary when collecting these sensitive data.
@@
The optimality of piecewise-based mechanisms remains a challenging problem. We will see that existing instantiations are heuristic forms of the $3$-piecewise mechanism (TPM). As a special case with $3$ pieces and pre-defined forms of pieces, TPM is far from enough to study the optimality of piecewise-based mechanisms.  From the evidence of the staircase Laplace mechanism (Geng and Viswanath, 2014) for unbounded numerical data, the asymptotically optimal mechanism has a staircase (multiple pieces) form. For categorical data, a staircase Randomized Response mechanism (SRR) (Wang et al., 2022) improves data utility in location collection compared to the general RR mechanism. Numerous indicators suggest that increasing the variety of probabilities in the data domain, i.e. using more pieces, can improve data utility.  In light of these examples, a fundamental question for piecewise-based mechanisms is: what is the optimal instantiation of piecewise-based mechanism? In the design of piecewise-based mechanisms, the number of pieces, their probabilities and sizes can be arbitrary. Finding the optimal instantiation within such a large design space is challenging, as it requires optimizing the number of pieces, their probabilities and sizes simultaneously.
@@
  • (Theoretical and experimental evaluations) We provide theoretical analyses of data utility and experimental evaluations on two common applications of LDP: distribution and mean estimation. The results prove and validate our mechanisms’ advantages over existing mechanisms. Codes are available at https://github.com/ZhengYeah/Optimal-GPM.
@@ 2.2. Local Differential Privacy
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function ($pdf$). Intuitively, $\varepsilon$-LDP represents the difficulty of distinguishing between $x_{1}$ and $x_{2}$ given $y$. A lower privacy parameter $\varepsilon\in[0,+\infty)$ means higher privacy.  For example, $\varepsilon=0$ requires $\mathcal{M}$ maps two arbitrary inputs to any output $y$ with the same probability, thus $y$ contains no distribution information about $x$, making any hypothesis-testing method to infer the sensitive $x$ powerless even with known $\mathcal{M}$.
@@ 2.3. Piecewise-based Mechanisms
When the input domain $\mathcal{D}$ is both continuous and bounded, the state-of-the-art mechanisms to achieve $\varepsilon$-LDP are piecewise-based mechanisms. We summarize these mechanisms as heuristic forms of the following definition.
@@
  • Not optimal in data utility. None of the existing instantiations provided closed forms for the optimal data utility. Meanwhile, they also assume invariable length $r_{x,\varepsilon}-l_{x,\varepsilon}$ of the central piece for all $x$, and symmetric probability $p_{\varepsilon}/\exp(\varepsilon)$ for the remaining two pieces. However, a general-form piecewise-based mechanism can have more pieces, unfixed piece lengths, and asymmetric probabilities, potentially improving data utility.
@@ 3.1. Framework for Deriving the Optimal GPM
  • (Solving difficulty) Formulation (2) is a min-max problem, and the integrand $\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}$ is non-linear; it is a non-convex problem whose global optimal hard to solve.
@@ 3.3. Closed Form for Classical Domain
The theoretical MSE allows us to analytically compare the optimal GPM with existing mechanisms. For mechanisms defined on $\mathcal{D}=[0,1)$, e.g. SW (Li et al., 2020), we can compare their MSE with the above result. For mechanisms defined on other domains, e.g. PM (Wang et al., 2019) on $\mathcal{D}=[-1,1)$, we can transform the optimal GPM to $[-1,1)$ and compare their MSE. Detailed comparison with them is presented in the evaluation section.
@@ 4.1. Reduced Forms
Figure 5. Reduced forms of solving the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$. Optimizations under $\mathcal{L}_{\text{mod}}$ can be reduced to those under $\mathcal{L}$.
@@ 4.2. Closed Form for Circular Domain
$\begin{split}&MSE(\mathcal{M}(x),x)=MSE(\mathcal{M}(\pi),\pi)\\ =&2\left(\int_{0}^{l_{\pi,\varepsilon}^{\textup{mod}}}(y-\pi)^{2}\frac{p_{\varepsilon}}{\exp(\varepsilon)}\mathrm{d}y+\int_{l_{\pi,\varepsilon}^{\textup{mod}}}^{\pi}(y-\pi)^{2}p_{\varepsilon}\mathrm{d}y\right).\end{split}$
@@
$MSE(\mathcal{M}(x),x)=\frac{2}{3}\left((\pi^{3}-C^{3})\frac{p_{\varepsilon}}{\exp(\varepsilon)}+C^{3}p_{\varepsilon}\right),$
@@
We have shown that the MSE at $x$ is the same as that at $\pi$. Then the calculation is straightforward by plugging in the values of $p_{\varepsilon}$ and $l_{\pi,\varepsilon}^{\textup{mod}}$. ∎
@@ 5.2. Mean Estimation
To estimate the mean of $\mathcal{X}$ from $\mathcal{Y}$, the collector uses the estimator $\hat{\mu}=\sum_{i=1}^{n}y_{i}/n$. The accuracy of this estimator is measured by $|\hat{\mu}-\mu|$, where $\mu$ is the true mean of $\mathcal{X}$. Our mechanisms guarantee the optimal error between each $x_{i}$ and $y_{i}$, which tends to give the smallest $|\hat{\mu}-\mu|$ for any $n$ among GPM.
Typically, a mean estimator may also need to be unbiased, i.e. $\mathrm{E}[\hat{\mu}]=\mu$. This constraint translates to $\mathrm{E}[y_{i}]=\mathrm{E}[\mathcal{M}(x_{i})]=x_{i}$.777 Actually, unbiasedness for numerical data is not as important as for categorical data, as it is for a single data point $x_{i}$. For numerical data with infinite possible data points, the probability of having a specific $x_{i}$ is zero. This is unachievable for same-domain mapping $\mathcal{M}:\mathcal{D}\to\mathcal{D}$ on classical domains, as the endpoints of $\mathcal{D}$ cannot be the mean value (or center) of any distribution over $\mathcal{D}$. For example, for any LDP mechanism $\mathcal{M}:[0,1)\to[0,1)$, distribution of $\mathcal{M}(0)$ can not be unbiased. So the mechanism in Theorem 3.7 is biased.
@@
Under the $3$-GPM, an unbiased mechanism $\mathcal{M}$ with a variable output domain $\tilde{\mathcal{D}}_{\varepsilon}$ can be analytically derived by incorporating the unbiasedness constraint. As a complement to Theorem 3.7, we provide Theorem 5.1 for mean estimation [^8] However, for GPM, unbiasedness does not always result in better data utility for mean estimation, as it comes at the cost of a larger output domain, which leads to higher errors (as shown in Figure 17). in the classical domain.
@@
Table 2 summarizes the optimal GPM for both distribution and mean estimation in three domain types. Theorems use a concrete domain $\mathcal{D}$ for the sake of clarity. We do not give closed-form optimal GPM for distribution estimation on $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ because $\tilde{\mathcal{D}}_{\varepsilon}$ can not be restricted by constraints as in mean estimation.
@@ 7. Evaluations
  • PM-C and SW-C: the compressed versions of PM and SW for $\mathcal{D}\to\mathcal{D}$. For the best potential of PM and SW, we adapt them to $\mathcal{D}\to\mathcal{D}$ as PM-C and SW-C by linearly compressing their output domain $\tilde{\mathcal{D}}_{\varepsilon}$ to $\mathcal{D}$, i.e. transformation invariants, which maintains the privacy level.
We also compare OGPM’s expected error with non-piecewise-based mechanisms that can be applied to bounded domains:
  • Variants of the Laplace mechanism: including the staircase mechanism (Geng and Viswanath, 2014), Laplace mechanism with post-processing by truncation (T-Laplace), and the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020), which designs a bounded Laplace-shape distribution.
@@ 7.1. Expected Errors
  • worst-case error: the largest $Err(x)$ value among the whole domain. PM (Wang et al., 2019) and PTT (Ma et al., 2024) use this error to evaluate data utility.
@@ 7.1.3. Comparison with the Staircase Mechanism, T-Laplace, and B-Laplace
In addition to piecewise-based mechanisms, the Laplace mechanism and its variants can also be applied to the classical domain to achieve LDP. Among these, the staircase mechanism (Geng and Viswanath, 2014) claims to be optimal under certain assumptions. For the input domain $\mathcal{D}=[0,1)$ (i.e. sensitivity $\Delta=1$) and error metric $\mathcal{L}=|y-x|$, its expected error is given by Theorem 3 in (Geng and Viswanath, 2014): $\exp(\varepsilon/2)/(\exp(\varepsilon)-1)$.  Another approach involves using the Laplace mechanism with truncation (Holohan et al., 2020), referred to here as T-Laplace for convenience. T-Laplace preserves the privacy guarantees of the Laplace mechanism while reducing the expected error, particularly for data points near the endpoints or for small $\varepsilon$ values. Additionally, the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020) introduces a redesigned bounded Laplace-shaped distribution tailored for bounded domains.999 Appendix B.8 provides details on the expected error of B-Laplace.
@@ 7.1.4. Comparison with the Purkayastha Mechanism
The paper “Differential Privacy for Directional Data” (Weggenmann and Kerschbaum, 2021) introduces two mechanisms for data on spheres $\mathbb{S}^{n-1}$: the VMF mechanism (ensuring indistinguishability of any two points with distance through the sphere) and the Purkayastha mechanism (ensuring indistinguishability of any two points with distance along the sphere). When $n=2$, the sphere $\mathbb{S}^{1}$ corresponds to a circle, making the Purkayastha mechanism a counterpart of OGPM in the circular domain. Therefore, we compare them in the circular domain.101010 We omit the comparison with the VMF mechanism also because (i) it has been shown that the Purkayastha mechanism outperforms the VMF mechanism (with the same sensitivity $\Delta_{\measuredangle}=\pi$ for sphere $\mathbb{S}^{1}$, e.g. Figure 5 and 10 in (Weggenmann and Kerschbaum, 2021)), and (ii) the expected error of the VMF mechanism lacks a closed-form expression (Theorem 17 in (Weggenmann and Kerschbaum, 2021)), making it complex to compute.
@@ 7.2.3. Mean Estimation
Figure 17 shows the comparison of the errors of mean estimation in the classical and circular domains. OGPM consistently outperforms other mechanisms in both types of domains, with significantly lower errors.  In the classical domain, we also compare with OGPM-U, which is designed for unbiased mean estimation, but it has an enlarged output domain, leading to a higher distance between the sensitive and randomized data. We can see that such larger output domain results in higher mean estimation errors when $\varepsilon$ is small, but OGPM-U outperforms SW when $\varepsilon\geq 5.5$.  In the circular domain, OGPM’s mean estimation error is negligible. We also observe that SW-C performs better than PM-C in this large-size domain.  Statistically, OGPM’s mean estimation error is $66.2\%$ of PM-C and $55.4\%$ of SW-C in the classical domain. In the circular domain, OGPM’s mean estimation error is merely $2.3\%$ of PM-C and $3.6\%$ of SW-C.

### v2

@@ Abstract.
This paper investigates the optimal design of piecewise-based mechanisms to maximize data utility under LDP. We demonstrate that existing piecewise-based mechanisms are heuristic instances of the $3$-piecewise mechanism, which is far from enough to study optimality. We generalize the $3$-piecewise mechanism to its most general form, i.e. $m$-piecewise mechanism with no pre-defined form of each piece. Under this form, we derive the closed-form optimal mechanism by combining analytical proofs and off-the-shelf optimization solvers. Next, we extend the generalized piecewise-based mechanism to the circular domain (along with the classical domain), defined on a cyclic range where the distance between the two endpoints is zero. By incorporating this property, we design the optimal mechanism for the circular domain, achieving significantly improved data utility compared with existing mechanisms.
@@ 1. Introduction
Numerical data with bounded domains is a fundamental data type in personal devices. These bounded domains can be categorized into two types: linear ranges, such as sensor readings in $[0,1)$, referred to as classical domain; and cyclic ranges, such as angular measurements in $[0,2\pi)$, referred to as circular domain. These types of data are crucial for third-party platforms to provide personalized services. However, collecting such data involves particular privacy concerns, as third-party collectors are potentially untrusted and the data often contain sensitive information. Simple anonymization techniques (Yang et al., 2013; Niu et al., 2014) have been proven insufficient to prevent privacy leakage (Wikipedia contributors, 2023; Narayanan and Shmatikov, 2008; Yu et al., 2007). Therefore, a provable privacy guarantee is necessary when collecting these sensitive data.
@@
The optimality of piecewise-based mechanisms remains a challenging problem. We will see that the existing instantiations are heuristic forms of the $3$-piecewise mechanism (TPM). As a special case with $3$ pieces and pre-defined forms of pieces, TPM is far from enough to study the optimality of piecewise-based mechanisms.  From the evidence of the staircase Laplace mechanism (Geng and Viswanath, 2014) for unbounded numerical data, the asymptotically optimal mechanism has a staircase (multiple pieces) form. For categorical data, a staircase Randomized Response mechanism (SRR) (Wang et al., 2022) improves data utility in location collection compared to the general RR mechanism. Numerous indicators suggest that increasing the variety of probabilities in the data domain, i.e. using more pieces, can improve data utility.  In light of these examples, a fundamental question for piecewise-based mechanisms is: what is the optimal instantiation of piecewise-based mechanism? In the design of piecewise-based mechanisms, the number of pieces, their probabilities and sizes can be arbitrary. Finding the optimal instantiation within such a large design space is challenging, as it requires optimizing the number of pieces, their probabilities and sizes simultaneously.
@@
  • (Theoretical and experimental evaluations) We provide theoretical analyses of data utility and experimental evaluations on two common applications of LDP: distribution and mean estimation. The results prove and validate our mechanisms’ advantages over existing mechanisms. The codes are available at https://github.com/ZhengYeah/Optimal-GPM.
@@ 2.2. Local Differential Privacy
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function ($pdf$). Intuitively, $\varepsilon$-LDP represents the difficulty of distinguishing between $x_{1}$ and $x_{2}$ given $y$. A lower privacy parameter $\varepsilon\in[0,+\infty)$ means higher privacy.  For example, $\varepsilon=0$ requires that $\mathcal{M}$ maps two arbitrary inputs to any output $y$ with the same probability, thus $y$ contains no distribution information about $x$, making any hypothesis-testing method to infer the sensitive $x$ powerless even with known $\mathcal{M}$.
@@ 2.3. Piecewise-based Mechanisms
When the input domain $\mathcal{D}$ is both continuous and bounded, the state-of-the-art mechanisms to achieve $\varepsilon$-LDP are piecewise-based mechanisms. We summarize these mechanisms as heuristic instances of the following definition.
@@
  • Not optimal in data utility. None of the existing instantiations provided closed forms for the optimal data utility. Meanwhile, they also assume an invariable length $r_{x,\varepsilon}-l_{x,\varepsilon}$ of the central piece for all $x$, and symmetric probability $p_{\varepsilon}/\exp(\varepsilon)$ for the remaining two pieces. However, a general-form piecewise-based mechanism can have more pieces, unfixed piece lengths, and asymmetric probabilities, potentially improving data utility.
@@ 3.1. Framework for Deriving the Optimal GPM
  • (Solving difficulty) Formulation (2) is a min-max problem, and the integrand $\mathcal{L}(y,x)\mathcal{P}_{\mathcal{M}(x)}$ is non-linear. It is a non-convex problem whose global optimal hard to solve.
@@ 3.3. Closed Form for Classical Domain
The theoretical MSE allows us to analytically compare the optimal GPM with existing mechanisms. For mechanisms defined on $\mathcal{D}=[0,1)$, e.g. SW (Li et al., 2020), we can compare their MSE with the above result. For mechanisms defined on other domains, e.g. PM (Wang et al., 2019) on $\mathcal{D}=[-1,1)$, we can transform the optimal GPM to $[-1,1)$ and compare their MSE. Detailed comparisons with them are presented in the evaluation section.
@@ 4.1. Reduced Forms
Figure 5. Reduced forms of solving the optimal $p_{i}$, $l_{i,x}^{\text{mod}}$ and $r_{i,x}^{\text{mod}}$. Optimizations under circular distance $\mathcal{L}_{\text{mod}}$ can be reduced to those under linear distance $\mathcal{L}$.
@@ 4.2. Closed Form for Circular Domain
$\begin{split}&\mathrm{MSE}[\mathcal{M}(x)]=\mathrm{MSE}[\mathcal{M}(\pi)]\\ =&2\left(\int_{0}^{l_{\pi,\varepsilon}^{\textup{mod}}}(y-\pi)^{2}\frac{p_{\varepsilon}}{\exp(\varepsilon)}\mathrm{d}y+\int_{l_{\pi,\varepsilon}^{\textup{mod}}}^{\pi}(y-\pi)^{2}p_{\varepsilon}\mathrm{d}y\right).\end{split}$
@@
$\mathrm{MSE}[\mathcal{M}(x)]=\frac{2}{3}\left((\pi^{3}-C^{3})\frac{p_{\varepsilon}}{\exp(\varepsilon)}+C^{3}p_{\varepsilon}\right),$
@@
We have shown that the MSE at $x$ is the same as that at $\pi$. Then the calculation is straightforward by plugging in the values of $p_{\varepsilon}$ and $l_{\pi,\varepsilon}^{\textup{mod}}$ in Theorem 4.2. ∎
@@ 5.2. Mean Estimation
To estimate the mean of $\mathcal{X}$ from $\mathcal{Y}$, the collector uses the estimator $\hat{\mu}=\sum_{i=1}^{n}y_{i}/n$. The accuracy of this estimator is measured by $|\hat{\mu}-\mu|$, where $\mu$ is the true mean of $\mathcal{X}$. Our mechanisms guarantee the optimal error between each $x_{i}$ and $y_{i}$, which in turn leads to the smallest $|\hat{\mu}-\mu|$ among all GPMs under this metric.
Typically, a mean estimator may also need to be unbiased, i.e. $\mathrm{E}[\hat{\mu}]=\mu$. This constraint translates to $\mathrm{E}[y_{i}]=\mathrm{E}[\mathcal{M}(x_{i})]=x_{i}$.777 Actually, unbiasedness for numerical data is not as important as for categorical data, as it is for a single data point $x_{i}$. When the dataset $\mathcal{X}$ is not concentrated around a single point, an unbiased mechanism may not necessarily provide better performance. This is unachievable for same-domain mapping $\mathcal{M}:\mathcal{D}\to\mathcal{D}$ on classical domains, as the endpoints of $\mathcal{D}$ cannot be the mean value (or center) of any distribution over $\mathcal{D}$. For example, for any LDP mechanism $\mathcal{M}:[0,1)\to[0,1)$, distribution of $\mathcal{M}(0)$ can not be unbiased. So the mechanism in Theorem 3.7 is biased.
@@
Hypothesis 5.1 is a natural extension of Hypothesis 3.7, as the output domain $\tilde{\mathcal{D}}_{\varepsilon}$ becomes explicit once $\varepsilon$ is specified. Under the $3$-GPM, an unbiased mechanism $\mathcal{M}$ with a variable output domain $\tilde{\mathcal{D}}_{\varepsilon}$ can be analytically derived by incorporating the unbiasedness constraint. As a complement to Theorem 3.7, we provide Theorem 5.1 for mean estimation in the classical domain.
@@
Table 2 summarizes the optimal GPM for both distribution and mean estimation in three domain types. Theorems use a concrete domain $\mathcal{D}$ for the sake of clarity. We do not give closed-form optimal GPM for distribution estimation on $\mathcal{D}\to\tilde{\mathcal{D}}_{\varepsilon}$ because $\tilde{\mathcal{D}}_{\varepsilon}$ can not be easily concretized by constraints as in mean estimation.
@@ 7. Evaluations
  • PM-C and SW-C: the compressed versions of PM and SW for $\mathcal{D}\to\mathcal{D}$. For the best potential of PM and SW, we adapt them to $\mathcal{D}\to\mathcal{D}$ as PM-C and SW-C by linearly compressing their output domain $\tilde{\mathcal{D}}_{\varepsilon}$ to $\mathcal{D}$, i.e. by transformation invariants, which maintains the privacy level.
We also compare OGPM’s expected error with non-piecewise-based mechanisms that can be applied to bounded numerical domains:
  • Variants of the Laplace mechanism: including the staircase mechanism (Geng and Viswanath, 2014), Laplace mechanism with post-processing by truncation (T-Laplace), and the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020), which redesigns a bounded Laplace-shape distribution.
@@ 7.1. Expected Errors
  • worst-case error: the largest $Err(x)$ value among the whole domain. PM (Wang et al., 2019) and PTT (Ma et al., 2024) also use this error to evaluate data utility.
@@ 7.1.3. Comparison with the Staircase Mechanism, T-Laplace, and B-Laplace
In addition to piecewise-based mechanisms, the Laplace mechanism and its variants can also be applied to the classical domain to achieve LDP. Among these, the staircase mechanism (Geng and Viswanath, 2014) claims to be optimal under certain assumptions. For the input domain $\mathcal{D}=[0,1)$ (i.e. sensitivity $\Delta=1$) and error metric $\mathcal{L}=|y-x|$, its expected error is given by Theorem 3 in (Geng and Viswanath, 2014): $\exp(\varepsilon/2)/(\exp(\varepsilon)-1)$.  Another approach involves using the Laplace mechanism with truncation (Holohan et al., 2020), referred to here as T-Laplace for convenience. T-Laplace preserves the privacy guarantees of the Laplace mechanism while reducing the expected error, particularly for data points near the endpoints or for small $\varepsilon$ values. Additionally, the bounded Laplace mechanism (B-Laplace) (Holohan et al., 2020) introduces a redesigned bounded Laplace-shaped distribution tailored for bounded domains.888 Appendix B.8 provides details on the expected error of B-Laplace.
@@ 7.1.4. Comparison with the Purkayastha Mechanism
The paper “Differential Privacy for Directional Data” (Weggenmann and Kerschbaum, 2021) introduces two mechanisms for data on spheres $\mathbb{S}^{n-1}$: the VMF mechanism (ensuring indistinguishability of any two points with distance through the sphere) and the Purkayastha mechanism (ensuring indistinguishability of any two points with distance along the sphere). When $n=2$, the sphere $\mathbb{S}^{1}$ corresponds to a circle, making the Purkayastha mechanism a counterpart of OGPM in the circular domain. Therefore, we compare them in the circular domain.999 We omit the comparison with the VMF mechanism also because (i) it has been shown that the Purkayastha mechanism outperforms the VMF mechanism (with the same sensitivity $\Delta_{\measuredangle}=\pi$ for sphere $\mathbb{S}^{1}$, e.g. Figure 5 and 10 in (Weggenmann and Kerschbaum, 2021)), and (ii) the expected error of the VMF mechanism lacks a closed-form expression (Theorem 17 in (Weggenmann and Kerschbaum, 2021)), making it complex to compute.
@@ 7.2.3. Mean Estimation
Figure 17 shows the comparison of the errors of mean estimation in the classical and circular domains. OGPM consistently outperforms other mechanisms in both types of domains, with significantly lower errors.  In the classical domain, we also compare with OGPM-U, which is specifically designed for unbiased mean estimation. Since the Accelerometer dataset is concentrated around zero, it particularly favors unbiased mechanisms, as their outputs tend to average closely to the true mean. We can see that OGPM-U achieves significantly lower errors than OGPM when $\varepsilon$ is small.  In the circular domain, OGPM is unbiased, having a negligible mean estimation error. We also observe that SW-C outperforms PM-C in this large-scale domain.  Statistically, OGPM’s mean estimation error is $66.2\%$ of PM-C and $55.4\%$ of SW-C in the classical domain. In the circular domain, OGPM’s mean estimation error is merely $2.3\%$ of PM-C and $3.6\%$ of SW-C.